OpenStudy (anonymous):

Help! Fan and medal: what is the slope intercept form of a line if the line goes through the points 65,65 and 74,74?

OpenStudy (anonymous):

I got to this answer but I'm pretty sure that's not right: M = 66-74/65-74 = -8/-9 = 8/9 y – y1 = 8/9(x – x1) y – 66 = 8/9(x – 65) y = 8/9 x -65 + 66 y = 8/9x + 1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i think you made a typo

OpenStudy (anonymous):

Where?

OpenStudy (misty1212):

\[\frac{74-65}{74-65}\] both x and y in the first one where 65 in the question

OpenStudy (anonymous):

oops the real question is 65,66

OpenStudy (misty1212):

oh

OpenStudy (freckles):

hint: remember a(b+c)=ab+ac not ab+c

OpenStudy (misty1212):

then your slope is right, but your y intercept is wrong

OpenStudy (misty1212):

damn that distributive law !!

OpenStudy (anonymous):

@freckles so it would be: M = 66-74/65-74 = -8/-9 = 8/9 y – y1 = 8/9(x – x1) y – 66 = 8/9(x – 65) y -66 = 8/9x -8/9(65) y = 8/9x -57.77 +66 y = 8/9x -57.77 +66 y = 8/9 - 8.23

OpenStudy (freckles):

\[(65,66) \text{ and } (74,74) \\ m=\frac{74-66}{74-65}=\frac{8}{9} \\ y-y_1=m(x-x_1) \\ y-66=\frac{8}{9}(x-65) \\ y-66=\frac{8}{9}x-\frac{8}{9} \cdot 65 \\ y-66=\frac{8}{9}x-\frac{520}{9} \\ y=\frac{8}{9}x-\frac{520}{9}+66 \\ y=\frac{8}{9}x-\frac{520}{9}+\frac{594}{9}\] \[y=\frac{8}{9}x+\frac{594-520}{9}\] You know if you are suppose to leave it in fractional form I will let you do 594-520 :)

OpenStudy (freckles):

fractional form also exact form since you sorta rounded 8*65/9

OpenStudy (anonymous):

Okay thanks so much

OpenStudy (freckles):

np